1. Consider the following reaction: 2NiO(s)→2Ni(s)+O2(g) If O2 is collected over water at 40.0 ∘C and a total pressure of 744 mmHg , what volume of gas will be collected for the complete reaction of 24.02 g of NiO?
Answer #1

Mass of NiO reacted = 24.02 g

Molar mass of NiO = 74.6928 g/mol

Moles of NiO reacted = mass/ molar mass = 24.02 g / ( 74.6928 g/mol) = 0.322 mol

From reaction

2 mol of NiO reacts to produce 1 mol of O2

So, moles of O2 produced (n) = 0.322 /2 = 0.161 mol

Temperature (T) = 40.0oC = 40 + 273.15 = 313.15 K

Pressure of O2 = Total pressure - Pressure of water at 40oC = 744 -55.5 =688.5 mmHg = 0.906 atm

[ 1 atm = 760 mmHg]

Gas constant (R) = 0.0821 L. atm / mol.K

From Ideal gas equation

PV = nRT

V = nRT/P = (0.161 mol * 0.0821 L. atm / mol.K * 313.15 K) / 0.906 atm = 4.22 L